# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

### v1

@@ Abstract
Understanding time is a pivotal aspect of human cognition, crucial in the broader framework of grasping the intricacies of the world. Previous studies typically focus on specific aspects of time, lacking a comprehensive temporal reasoning benchmark. To address this issue, we propose TimeBench, a comprehensive hierarchical temporal reasoning benchmark that covers a broad spectrum of temporal reasoning phenomena, which provides a thorough evaluation for investigating the temporal reasoning capabilities of large language models. We conduct extensive experiments on popular LLMs, such as GPT-4, LLaMA2, and Mistral, incorporating chain-of-thought prompting. Our experimental results indicate a significant performance gap between the state-of-the-art LLMs and humans, highlighting that there is still a considerable distance to cover in temporal reasoning. We aspire for TimeBench to serve as a comprehensive benchmark, fostering research in temporal reasoning for LLMs. Our resource is available at https://github.com/zchuz/TimeBench
@@ 1 Introduction
Time flies over us, but leaves its shadow behind. Time continually moves forward, threading through the fabric of people’s lives. Understanding time is a crucial part of human comprehension of the world. Imagine witnessing the blossoming of flowers, and you’ll connect it to the arrival of spring. The ponder within it encompasses the intricate interplay of world knowledge, causality, and event temporal relationships. Temporal reasoning, in contrast to reasoning of a singular nature, comes with inherent complexity, encompassing implicit arithmetic, logical implications, and world knowledge. It is a form of integrated reasoning built upon foundational reasoning like mathematical and logical reasoning (Cobbe et al., 2021; Mishra et al., 2022; Yu et al., 2020). Recently, large language models (LLM) have demonstrated remarkable performance in complex reasoning (Hendrycks et al., 2021; Srivastava et al., 2022; Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023; Touvron et al., 2023), but their performance in intricate temporal reasoning still lacks a definitive conclusion. The research community currently lacks a comprehensive benchmark, which makes it challenging to quantify the temporal reasoning capabilities of LLMs.
@@
Recent benchmarks for temporal reasoning typically focus only on a few aspects, such as temporal commonsense or temporal question answering (Zhou et al., 2019; Chen et al., 2021; Dhingra et al., 2022; Wang and Zhao, 2023). Due to the inherent complexity of temporal reasoning, it is challenging to accurately measure models’ temporal reasoning capabilities based on limited aspects.
@@
To address this issue, we propose TimeBench, a comprehensive and hierarchical temporal reasoning benchmark, which is aligned with intrinsic real-world scenarios. Specifically, drawing inspiration from the human cognitive process of transitioning from abstraction and concreteness to integration (Barsalou et al., 2018), we categorize temporal reasoning into three levels: symbolic temporal reasoning, commonsense temporal reasoning, and event temporal reasoning. These levels respectively represent understanding abstract time expression, grasping concrete world knowledge, and integrating and applying this knowledge in real-world scenarios. It comprises 10 tasks with 16 sub-tasks, covering a broad spectrum of temporal reasoning phenomena. Besides, prior work typically featured only a singular task form, overly simplistic and insufficient in capturing the model’s performance in real-world scenarios. In contrast to them, we incorporate four distinct task forms, offering a more realistic simulation of challenges models encounter in real-world scenarios.
@@
To quantify the temporal reasoning capabilities of contemporary LLMs, we extensively assess widely-used LLMs, including closed-source models such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), as well as open-source like LLaMA2 (Touvron et al., 2023), Vicuna-1.5 (Chiang et al., 2023), Mistral (Jiang et al., 2023), Baichuan2 (Yang et al., 2023), ChatGLM3 Zeng et al. (2023) and FLAN-T5 (Shen et al., 2023). We conduct experiments under zero-shot and few-shot settings, combining commonly used reasoning techniques, chain-of-thought prompting (Kojima et al., 2022; Wei et al., 2022; Chu et al., 2023a). The experimental results suggest that GPT-4 continues to outperform other models, showcasing strong temporal reasoning capabilities, as shown in Figure 1. Nevertheless, there is still a considerable gap with humans. On the contrary, open-source models generally exhibit subpar performance, lagging significantly behind GPT-3.5 and showing notable shortcomings in commonsense reasoning. These findings indicate that there is still significant room for improvement in models’ temporal reasoning capabilities. In addition, we also observe that, unlike tasks such as mathematical or logical reasoning, chain-of-thought prompting does not consistently enhance model performance; in some cases, it might even impair performance.
We summarize our contributions as follows: (1) We introduce TimeBench, a comprehensive and hierarchical temporal reasoning benchmark. (2) We conduct extensive experiments, revealing a significant gap between even sota LLM and humans, indicating substantial research opportunities in this field. (3) By conducting a thorough analysis, we uncover the dilemmas that models face in temporal reasoning and identify potential solutions.
Figure 2: Distribution of category, task and subtask in TimeBench. Degree of arc indicates the ratio of datas.
@@ 2.1 Benchmark Design Principal
TimeBench focuses on a comprehensive evaluation of the temporal reasoning capabilities of large language models in challenging and complex scenarios. To achieve this goal, we summarize the difficulties and challenges faced in temporal reasoning, categorize it into three levels, and meticulously design tasks that better align with complex real-world scenarios.
Just as the human cognitive process unfolds from foundational cognition and conceptual understanding to practical reasoning, we delineate temporal reasoning into three hierarchical levels. Specifically, TimeBench categorizes temporal reasoning into symbolic, commonsense and event temporal reasoning, covering 10 datasets with a total of 16 subtasks. (1) Symbolic Temporal Reasoning focuses on comprehension of fundamental abstract temporal expressions. (2) Temporal Commonsense Reasoning emphasizes the mastery of temporal principles, concepts and world knowledge. (3) Event Temporal Reasoning concentrates on modeling the temporal relationships of events within authentic scenarios. Furthermore, to better align with real-world scenarios, we employ diverse question formats.
@@ 2.2 Difficulties and Challenges
We outline the necessary capabilities and challenges encountered from a human cognition perspective during temporal reasoning, and language models confront similar obstacles. Detailed information on TimeBench and challenges involved in each task is presented in Table 6.
TimeX Understanding
Time expressions (TimeX) denote words or phrases that convey information about time and represent the simplest and most basic units of expressing time, such as in August 2008, which is fundamental in comprehending time.
@@
assesses the understanding of temporal aspects of world knowledge, including event order, event duration, event typical time, and event frequency, which is crucial for language models to comprehend everyday scenarios.
@@
evaluates the grounding ability to establish correspondence between events and time, enabling the model to comprehend the development and changes of events as they unfold over time.
@@
not only involve event-time grounding but also introduce multi-hop relative connections. Models with this capability can better handle temporal reasoning in complex scenarios involving multiple events.
@@
2.3 Symbolic Temporal Reasoning
To evaluate the language model’s comprehension of abstract time expressions, we utilize two symbolic reasoning tasks stripped of semantic content: TimeX arithmetic and TimeX inference. Table 1 shows examples of symbolic temporal reasoning.
TimeX Arithmetic (Tan et al., 2023)
assesses the model’s grasp of abstract date calculation. When provided with a date, the model needs to accurately calculate the date a certain amount of time before or after the given date.
TimeX NLI (Thukral et al., 2021)
focuses on the logical entailment relationships among abstract TimeX, including three aspects: order (s1), duration (s2), and duration with unit conversion (s3).
MCTACO
@@ 2.3 Symbolic Temporal Reasoning
2.4 Commonsense Temporal Reasoning
We measure the model’s mastery of temporal common and world knowledge, along with its capacity for reasoning based on these insights. Table 2 presents examples of temporal commonsense reasoning in QA and generation forms.
MCTACO (Zhou et al., 2019)
evaluates diverse commonsense knowledge from different aspects of events, including duration, frequency, order, stationary and typical event time.
DurationQA (Virgo et al., 2022)
focuses specifically on temporal commonsense reasoning in the spectrum of event duration.
TimeDial (Qin et al., 2021)
considers temporal commonsense reasoning in dialogue scenarios and involves various aspects of commonsense associated with duration, order, and world knowledge.
SituatedGen (Zhang and Wan, 2023)
considers generative commonsense reasoning in a constrained text generation scenario. Given a set of contrasting keywords, the model needs to choose appropriate keywords for each sentence and generate a pair of contrasting sentences that satisfy temporal commonsense.
TimeQA
C: … He worked in Utrecht for the firm of P Smits & de Wolf from 1864 to 1867 and then returned to Widnes …
@@ 2.4 Commonsense Temporal Reasoning
Table 3: Examples of event temporal reasoning. C=context, Q=question, A=answer
@@ 2.5 Event Temporal Reasoning
TimeQA (Chen et al., 2021)
requires the model to answer time-sensitive questions based on context containing numerous time-involved facts. It is categorized into explicit reasoning and implicit reasoning based on time indicators (before, in, etc.).
MenatQA (Wei et al., 2023)
introduces temporal factors to elicit implicit temporal reasoning, including time scope change, disruption of facts, and counterfactual questions, which provides a more in-depth assessment of implicit reasoning ability on event-time relations.
TempReason (Tan et al., 2023)
removes irrelevant context and focuses on implicit temporal reasoning within structured facts, investigating the model’s capability boundaries. It involves event-time reasoning and event-event reasoning.
TRACIE Zhou et al. (2021)
evaluates the model’s comprehension of temporal order between implicit events. The model needs to identify events implied in the context and then determine their chronological order.
2.6 Task Format and Evaluation Metrics
Real-world scenarios are diverse in task forms. To investigate the model’s authentic performance, we select diverse task forms, covering a wide spectrum of temporal reasoning phenomena.
Multi-choice & Multi-answer Questions
MCQ tasks require models to select the only correct answer from the options. However, this task form has shortcuts and may not truly reflect the model’s capability. To address this, we use the MCMAQ form, where the model needs to select all possible correct answers from options. In our task, each question includes four options, with two of them being correct. Task: MCTACO, DurationQA, TimeDial.
Natural Language Inference
is the task of determining the logical relationship between two pieces of text. Specifically, given a premise and a hypothesis, the model needs to determine whether the hypothesis can be inferred from the premise and output entailment, contradiction, or neutral. Our tasks focus on the entailment in temporal domains. Tasks: TimeX-NLI, TRACIE.
Free-form Reading Comprehension
requires models to answer questions based on the provided context, and the answer is free-form without pre-defined restrictions. Tasks: TimeQA, MenatQA, TempReason, Date Calculation.
Constrained Text Generation
refers to the task of generating text under certain constraints. Our task is keyword-constrained text generation, where the model takes keywords as input and outputs sentences that include those keywords. Task: SituatedGen.
Metrics
We adopt the evaluation metrics from previous work. Accuracy is used for NLI and date calculation tasks. MAMCQ tasks are measured using option-based EM and F1. FRC tasks (w/o date calc.) are evaluated with token-based EM and F1. For CTG tasks, we use a composite generation metrics, with more details in Appendix A.2.
@@ 3 Methodology
We perform evaluations using the prompt-based approach, including standard I-O prompting and chain-of-thought prompting. Experiments are conducted under both zero-shot and few-shot settings.
Standard I-O Prompting
@@
$\mathrm{prompt_{zs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q}\}$ (1)
$\mathrm{prompt_{fs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q_{1}}\}\{\mathrm{A_{1}}\}..\{\mathrm{Q}\}$ (2)
@@
We use identical instructions to standard input-output prompting. In the zero-shot chain-of-thought scenario, following Zeroshot CoT (Kojima et al., 2022), we add a reasoning trigger Let’s think step by step after questions to perform chain-of-thought reasoning. In the few-shot chain-of-thought scenario, we manually craft rationales for each task as demonstrations to guide the model to reason step-by-step. The instructions and demonstrations can be found in the Appendix A.4.

### v2

@@ Abstract
Grasping the concept of time is a fundamental facet of human cognition, indispensable for truly comprehending the intricacies of the world. Previous studies typically focus on specific aspects of time, lacking a comprehensive temporal reasoning benchmark. To address this, we propose TimeBench, a comprehensive hierarchical temporal reasoning benchmark that covers a broad spectrum of temporal reasoning phenomena. TimeBench provides a thorough evaluation for investigating the temporal reasoning capabilities of large language models. We conduct extensive experiments on GPT-4, LLaMA2, and other popular LLMs under various settings. Our experimental results indicate a significant performance gap between the state-of-the-art LLMs and humans, highlighting that there is still a considerable distance to cover in temporal reasoning. Besides, LLMs exhibit capability discrepancies across different reasoning categories. Furthermore, we thoroughly analyze the impact of multiple aspects on temporal reasoning and emphasize the associated challenges. We aspire for TimeBench to serve as a comprehensive benchmark, fostering research in temporal reasoning111 Data is available at: GitHub.
[…]
TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models
Zheng Chu1, Jingchang Chen1, Qianglong Chen2, Weijiang Yu3, Haotian Wang1, Ming Liu1,4thanks:  Corresponding Author., Bing Qin1,4 1Harbin Institute of Technology, Harbin, China 2Zhejiang University   3Sun Yat-sen University   4Peng Cheng Laboratory {zchu,jcchen,mliu,qinb}@ir.hit.edu.cn {chenqianglong.ai, wanght1998, weijiangyu8}@gmail.com
@@ 1 Introduction
Time flies over us, but leaves its shadow behind. Understanding time is a crucial part of human comprehension of the world. Envision the blossoming of flowers, and you’ll associate it with the arrival of spring. The ponder within it encompasses the intricate interplay of world knowledge, causality, and event temporal relationships. Temporal reasoning, in contrast to reasoning of a singular nature, comes with inherent complexity, encompassing implicit arithmetic, logical implications, and world knowledge. It is a form of integrated reasoning built upon foundational reasoning like mathematical and logical reasoning (Cobbe et al., 2021; Mishra et al., 2022; Yu et al., 2020). Recently, large language models (LLMs) have demonstrated remarkable performance in complex reasoning (Hendrycks et al., 2021; Srivastava et al., 2022; Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023; Touvron et al., 2023), but their performance in temporal reasoning has not yet been extensively explored.
@@
Recent research for temporal reasoning typically focuses only on a few aspects, such as temporal commonsense or temporal question answering (Zhou et al., 2019; Chen et al., 2021; Dhingra et al., 2022; Wang and Zhao, 2023). Due to the inherent complexity of temporal reasoning, it is challenging to accurately measure models’ temporal reasoning capabilities based on limited aspects.
@@
To address this issue, we propose TimeBench, a comprehensive and hierarchical temporal reasoning benchmark. Specifically, drawing inspiration from the human cognitive process of transitioning from abstraction and concreteness to integration (Barsalou et al., 2018), we categorize temporal reasoning into three levels: symbolic temporal reasoning, commonsense temporal reasoning, and event temporal reasoning. These levels respectively represent understanding abstract time expression, grasping concrete world knowledge, and integrating and applying this knowledge in real-world scenarios. TimeBench comprises 10 tasks with 16 sub-tasks, covering a broad spectrum of temporal reasoning phenomena. Besides, prior work typically features only a single task form, too simplistic to capture the model’s performance. In contrast, we incorporate four distinct task forms, offering a more realistic simulation of challenges.
@@
To quantify the temporal reasoning capabilities of contemporary LLMs, we extensively assess widely-used LLMs, including proprietary models such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), as well as open-source like LLaMA2 (Touvron et al., 2023), Vicuna-1.5 (Chiang et al., 2023), Mistral (Jiang et al., 2023), Baichuan2 (Yang et al., 2023a), ChatGLM3 Zeng et al. (2023) and FLAN-T5 (Chung et al., 2022). We conduct experiments under zero-shot and few-shot settings, combining commonly used reasoning techniques, chain-of-thought prompting (Kojima et al., 2022; Wei et al., 2022). The experimental results suggest that GPT-4 outperforms other models, showcasing strong temporal reasoning capabilities, as shown in Figure 1. Nevertheless, there is still a considerable gap between the strongest models and humans. On the contrary, open-source models show inferior performance in temporal reasoning, attributed to shortcomings in abstract time understanding, temporal relations modeling, and a lack of temporal commonsense. In addition, we also observe that chain-of-thought prompting does not yield a consistent improvement in performance. These findings indicate that there is still significant room for improvement in models’ temporal reasoning capabilities. Moreover, we have conducted a thorough analysis of the deficiencies and obstacles faced by models in temporal reasoning.
We aspire for temporal reasoning to garner increased attention within the research community. Our contributions can be summarized as follows:
  • We introduce TimeBench, a comprehensive and hierarchical benchmark to quantify the temporal reasoning abilities of LLMs.
  • We conduct extensive experiments with several LLMs, revealing a significant gap between even SOTA LLM and humans, indicating substantial research opportunities in this field.
  • By conducting a thorough analysis, we reveal the dilemmas that LLMs face in temporal reasoning and identify potential solutions.
@@ 2.1 Benchmark Design Principal
TimeBench focuses on a comprehensive evaluation of the temporal reasoning capabilities of large language models in challenging and complex scenarios. To achieve this goal, we summarize the difficulties and challenges faced in temporal reasoning, categorize them into three levels, and integrate diverse task formats to better align with the intricate nature of temporal reasoning.
Just as the human cognitive process unfolds from foundational cognition and conceptual understanding to practical reasoning, we delineate temporal reasoning into three hierarchical levels. Specifically, TimeBench categorizes temporal reasoning into symbolic, commonsense and event temporal reasoning, covering 10 datasets with a total of 16 subtasks. (1) Symbolic Temporal Reasoning focuses on the comprehension of fundamental abstract temporal expressions. (2) Temporal Commonsense Reasoning emphasizes the mastery of temporal principles, concepts and world knowledge. (3) Event Temporal Reasoning concentrates on modeling the temporal relationships between events and times within authentic scenarios.
@@ 2.2 Difficulties and Challenges
We delineate the essential competencies and the challenges that arise from a human cognitive standpoint in the realm of temporal reasoning, and language models confront similar challenges. We present the dataset statistics, task formats, and the associated challenges in Table 7.
Time Expression Understanding
Time expressions (TimeX) denote words or phrases that convey information about time and represent the simplest and most basic units of expressing time, such as in April 2000, after 2008. Grasping time expressions is the most foundational step in understanding temporal elements within the textual modality.
@@
assesses the understanding of temporal world knowledge, including event order, event duration, typical time, event frequency and stationary, which is crucial for language models to comprehend daily scenarios.
@@
assesses the model’s grounding capability to establish temporal relationships between events and their temporal context, thereby enabling models to grasp the progression and transformations of events as they dynamically evolve through time.
@@
not only involve event-time grounding but also introduce multi-hop relative connections between events. Models with this capability can better handle temporal reasoning in complex scenarios involving multiple events.
@@
2.3 Symbolic Temporal Reasoning
To evaluate the language model’s comprehension of abstract time expressions, we utilize two symbolic reasoning tasks stripped of semantic content: date arithmetic and time expression inference. Table 1 shows examples of symbolic temporal reasoning.
Date Arithmetic
(Tan et al., 2023) assesses the model’s grasp of abstract date calculation. When provided with a date, the model needs to accurately calculate the date a certain amount of time before or after the given date. The smallest unit is one day.
TimeX NLI
(Thukral et al., 2021) focuses on the logical entailment relationships among abstract TimeX, including three aspects: order (s1), duration (s2), and duration with unit conversion (s3).
C: … He worked in Utrecht for the firm of P Smits & de Wolf from 1864 to 1867 and then returned to …
@@ 2.3 Symbolic Temporal Reasoning
TempReason
Table 3: Examples of event temporal reasoning.
2.4 Commonsense Temporal Reasoning
We measure the model’s mastery of temporal commonsense and world knowledge, along with its capacity for reasoning based on these insights. Table 2 presents examples of temporal commonsense reasoning in QA and generation forms.
MCTACO
(Zhou et al., 2019) evaluates diverse commonsense knowledge from different aspects of events, including duration, frequency, order, stationary and typical event time.
DurationQA
(Virgo et al., 2022) focuses specifically on temporal commonsense reasoning in the spectrum of event duration.
TimeDial
(Qin et al., 2021) considers temporal commonsense reasoning in dialogue scenarios and involves various aspects of commonsense associated with duration, order, and world knowledge.
SituatedGen
(Zhang and Wan, 2023) considers generative commonsense reasoning in a constrained text generation scenario. Given a set of contrasting keywords, the model needs to choose appropriate keywords for each sentence and generate a pair of contrasting sentences that satisfy temporal commonsense.
@@ 2.5 Event Temporal Reasoning
TimeQA
(Chen et al., 2021) requires the model to answer time-sensitive questions based on context containing numerous time-involved facts. It is categorized into explicit reasoning and implicit reasoning based on time indicators (before, in, etc.).
MenatQA
(Wei et al., 2023) introduces time-sensitive factors to elicit implicit temporal reasoning, including time scope change, disruption of facts, and counterfactual questions, which provides a more in-depth assessment of implicit reasoning ability on event-time relations.
TempReason
(Tan et al., 2023) removes irrelevant context and focuses on implicit temporal reasoning within structured facts, investigating the model’s capability boundaries. It involves event-time reasoning and event-event reasoning.
TRACIE
(Zhou et al., 2021) evaluates the model’s comprehension of temporal order between implicit events. The model needs to identify events implied in the context and then determine their chronological order.
2.6 Task Formats and Evaluation Metrics
TimeBench is a multispectral benchmark encompassing four task types: free-form reading comprehension, natural language inference, constrained text generation, and multi-select questions. For detailed task types and their corresponding evaluation metrics, please refer to Appendix A.3 and A.4.
@@ 3 Methodology
We perform evaluations using the prompt-based approach, including standard prompting and chain-of-thought prompting. Experiments are conducted under both zero-shot and few-shot settings.
Standard Prompting
@@
$\mathrm{prompt_{zs}^{sp}}=\{\mathrm{INST}\}\{\mathrm{Q}\}$ (1)
$\mathrm{prompt_{fs}^{sp}}=\{\mathrm{INST}\}\{\mathrm{Q_{1}}\}\{\mathrm{A_{1}}\}..\{\mathrm{Q}\}$ (2)
@@
The instructions of CoT are the same as standard prompting. In the zero-shot setting, following Zeroshot CoT (Kojima et al., 2022), we add a reasoning trigger Let’s think step by step after questions to perform chain-of-thought reasoning. In the few-shot setting, we manually annotate CoT demonstrations for each task to guide the step-by-step reasoning. Prompts can be found in Appendix B.3.
@@ 4 Experimental Setup
Method Symbolic Commonsense Event Temporal Overall
TimeXNLI Arith DQA McT. TiD. SitGen TimeQA MenatQA TempR TRACIE Sym. Comm. Event Avg.
s1 s2 s3 Exp. Imp. Sco. Ord. Ctf. L2 L3
Human 98.0 96.0 92.0 100.0 80.8 87.1 97.8 100.0 93.3 91.1 85.6 87.3 79.9 97.1 95.3 82.5 96.5 91.4 89.0 91.5
GPT-4 85.3 73.3 53.3 100.0 64.8 88.3 94.6 88.6 73.7 51.0 72.4 54.8 28.7 92.4 95.9 62.8 78.0 84.1 66.5 73.7
+ FS CoT 92.0 84.0 64.0 100.0 55.1 72.3 93.4 - 66.9 52.8 65.3 52.6 25.9 96.9 94.6 66.4 85.0 73.6 65.2 72.1
GPT-3.5 52.0 68.4 31.6 63.6 67.7 71.2 76.4 79.1 66.1 48.4 43.2 51.6 17.9 84.7 78.0 55.0 53.9 73.6 55.6 59.7
+ FS CoT 51.6 71.8 36.6 84.4 41.2 38.1 71.1 - 68.0 47.0 42.5 41.7 37.8 89.9 76.6 50.2 61.1 50.1 56.7 56.6
[…]
+ FS CoT 52.0 73.0 39.0 79.5 62.3 79.1 61.1 - 64.3 43.0 57.7 45.2 53.1 87.5 81.6 67.0 60.9 67.5 62.4 63.0
[…]
+ FS CoT 40.0 61.0 37.0 52.0 59.3 68.8 40.8 - 59.4 49.1 58.4 43.8 44.1 78.0 68.2 58.0 47.5 56.3 57.4 54.5
[…]
+ FS CoT 37.0 52.0 36.0 25.5 56.9 67.0 41.9 - 45.6 36.1 50.9 38.0 57.3 59.7 57.7 50.0 37.6 55.3 49.4 47.4
[…]
+ FS CoT 50.0 56.0 34.0 47.0 62.0 69.3 43.8 - 58.2 49.6 49.8 40.1 45.6 81.3 65.6 60.0 46.8 58.4 56.3 54.2
[…]
+ FS CoT 30.0 56.0 34.0 34.0 57.0 69.5 44.5 - 51.2 40.7 46.4 32.6 46.3 61.5 64.1 53.0 38.5 57.0 49.5 48.1
[…]
+ FS CoT 57.0 63.0 35.0 54.0 61.8 45.7 57.3 - 60.4 46.2 57.2 47.9 33.2 65.9 67.9 57.0 52.3 54.9 54.5 54.0
[…]
+ FS CoT 47.0 68.0 32.0 46.0 53.9 64.3 56.5 - 52.5 24.5 35.0 40.2 22.5 79.4 60.3 54.0 48.3 58.2 46.1 49.1
Table 4: Experimental results under few-shot settings (standard prompting by default). denotes the base model without alignment. Global top-3 results are bold. Figure 8 provides a horizontal comparison of the performance of all models. Full results in Appendix B.2.
Method Symbolic Commonsense Event Temporal Overall
TimeXNLI Arith DQA McT. TiD. SitGen TimeQA MenatQA TempR TRACIE Sym. Comm. Event Avg.
s1 s2 s3 Exp. Imp. Sco. Ord. Ctf. L2 L3
Human 98.0 96.0 92.0 100.0 80.8 87.1 97.8 100.0 93.3 91.1 85.6 87.3 79.9 97.1 95.3 82.5 96.5 91.4 89.0 91.5
GPT-4 78.6 76.0 50.7 98.0 59.2 80.0 91.1 59.3 60.6 46.5 57.0 57.0 23.1 95.3 95.0 64.8 75.8 72.4 62.4 68.3
+ CoT 80.0 76.0 60.0 92.0 58.1 82.6 89.3 - 61.3 41.2 54.6 59.6 22.6 97.0 94.5 58.0 77.0 76.7 61.1 68.5
GPT-3.5 45.4 67.6 31.2 97.0 50.5 68.6 69.1 62.3 70.8 35.4 40.9 43.9 22.9 81.2 73.8 57.4 60.3 62.6 53.3 57.4
+ CoT 33.6 64.8 33.6 71.0 23.2 45.1 67.0 - 64.4 35.1 39.7 42.9 26.3 57.6 68.1 52.0 50.8 45.1 48.3 48.3
LLaMA270b 44.0 47.0 32.0 78.5 59.2 68.9 57.0 25.0 40.8 40.6 18.9 16.6 12.0 63.5 54.5 48.0 50.4 52.5 36.8 44.1
+ CoT 30.0 66.0 28.0 53.5 57.3 67.1 58.6 31.4 19.5 12.2 12.7 20.8 37.5 40.5 51.0 44.4 61.0 28.2 39.1
LLaMA213b 30.0 49.0 34.0 22.5 38.5 40.6 35.4 57.9 61.9 30.5 46.1 36.1 26.9 53.1 69.4 49.0 33.9 43.1 46.6 42.6
+ CoT 36.0 50.0 38.0 6.0 39.2 51.7 36.9 - 58.7 38.9 40.9 32.5 33.6 58.0 68.4 47.0 32.5 42.6 47.3 42.4
LLaMA27b 39.0 53.0 30.0 13.0 39.3 41.0 6.3 24.5 49.0 29.0 26.8 21.1 16.0 63.9 47.9 49.0 33.8 27.8 37.8 34.3
+ CoT 44.0 50.0 33.0 5.0 35.0 40.0 1.7 - 49.9 31.6 31.4 24.5 17.8 56.9 48.1 46.0 33.0 25.6 38.3 34.3
Baichuan213b 41.0 61.0 37.0 12.5 52.0 63.4 57.7 52.2 55.4 34.6 48.8 44.3 39.5 57.4 61.4 49.0 37.9 56.3 48.8 48.0
+ CoT 40.0 57.0 31.0 10.0 44.6 61.9 58.1 - 41.5 40.9 52.0 38.5 43.2 62.8 64.3 55.0 34.5 54.9 49.8 46.7
Baichuan27b 35.0 50.0 37.0 4.5 47.9 55.3 54.3 42.0 41.5 34.7 35.2 31.2 20.4 43.4 47.7 55.0 31.6 49.9 38.6 39.7
+ CoT 38.0 43.0 32.0 1.0 37.9 58.0 44.2 - 53.5 38.8 39.9 33.2 29.3 41.2 47.2 54.0 28.5 46.7 42.1 39.4
Vicuna1.513b 35.0 50.0 36.0 15.0 39.2 59.1 34.2 51.8 60.4 37.0 46.8 37.4 23.2 42.1 43.6 46.0 34.0 46.1 42.1 41.1
+ CoT 42.0 51.0 37.0 3.0 29.8 50.0 33.7 - 56.9 36.4 38.2 37.7 20.4 49.0 49.1 51.0 33.3 37.8 42.3 39.0
Vicuna1.57b 37.0 58.0 43.0 5.0 40.4 52.5 32.0 47.8 47.1 18.5 35.7 25.7 17.3 33.0 46.8 54.0 35.8 43.2 34.8 37.1
+ CoT 36.0 50.0 36.0 1.5 39.4 49.2 36.2 - 40.9 24.6 26.2 28.5 25.0 27.7 40.3 54.0 30.9 41.6 33.4 34.4
FLANT511b 53.0 63.0 43.0 0.0 52.0 65.0 47.7 49.5 61.7 26.8 33.6 52.2 21.8 87.9 83.9 64.0 39.8 53.6 54.0 50.3
+ CoT 56.0 66.0 45.0 0.0 49.7 63.4 42.7 - 64.4 28.2 41.6 50.2 30.6 79.5 68.9 55.0 41.8 51.9 52.3 49.4
Mistral7b 47.0 50.0 43.0 26.5 49.8 58.8 23.2 58.3 28.2 21.4 24.3 22.3 21.7 39.6 31.6 51.0 41.6 47.5 30.0 37.3
+ CoT 38.0 56.0 35.0 16.5 36.6 49.3 19.3 - 31.3 22.4 21.1 24.9 25.6 34.0 31.2 61.0 36.4 35.1 31.4 33.5
ChatGLM36b 38.0 50.0 34.0 2.0 34.1 43.6 56.7 38.9 41.2 31.7 33.8 26.0 32.2 57.0 54.0 50.0 31.0 43.3 40.7 39.0
+ CoT 27.0 49.0 37.0 0.0 24.8 37.1 44.8 - 41.7 25.4 34.6 28.1 41.2 44.5 52.0 48.0 28.3 35.6 39.4 35.7
Table 5: Experimental results under zero-shot settings (standart prompting by default). All models are alignment models (-chat or -instruct). Global top-3 results are bold.